OpenStudy (anonymous):

Plzzz help! Find the value of C(7,0) + C(7,1) + C(7,1) + C(7,2) +...+ C(7, 6) + (7, 6)

OpenStudy (ikram002p):

C(7,0) + C(7,1) + C(7,1) + C(7,2) +...+ C(7, 6) + (7, 6) ?? define C is it constant ??

OpenStudy (anonymous):

What exactly is the pattern?

OpenStudy (anonymous):

Is it \[ \sum_{k=0}^n2\binom {n}{k} \]?

OpenStudy (anonymous):

Because that is just \[ \sum_{k=0}^n2\binom {n}{k}= 2\sum_{k=0}^n\binom {n}{k}=2\times 2^n=2\times 2^7=256 \]

OpenStudy (anonymous):

But somehow you're missing terms.

OpenStudy (anonymous):

I mean to be really picky what you have is: \[ \left[\sum_{k=0}^n2\binom {n}{k}\right]-\binom 70 -2\binom 77=\left[\sum_{k=0}^n2\binom {n}{k}\right]-3=256-3=253 \]

OpenStudy (ikram002p):

mmmm the easy way ! C(7,0) + C(7,1) + C(7,1) + C(7,2) + C(7,3)+ C(7,4)+ C(7,5)+ C(7, 6) + (7, 6) \[=(C(7*n+2),\frac{ n(n+1) }{ 2 } +6)\] such that n=7.

OpenStudy (anonymous):

thank u all but i think i am getting even more confused Do u know this formula: \[^{n} C_{r} =\frac{ n! }{ r!(n-r)! }\] I think we are supposed to use this formula to solve this question. I have been trying soo hard but still I am not getting to anything almost 5 or 6 pages full of rough work I will try solving using the hint by @ikram002p so thanx @ikram002p

OpenStudy (anonymous):

i m sorry but i completely do not understand using your method @wio i knw u r using summation to solve but i do not knw tht method so thts why but ill give it a try as well so thnx to you @wio as well :)

OpenStudy (anonymous):

My method comes from the binomial theorem: \[ \sum_{k=0}^{n}\binom nk a^kb^{n-k}=(a+b)^n \]When you let \(a=1,b=1\) then: \[ \sum_{k=0}^{n}\binom nk 1^k1^{n-k}=(1+1)^n \]Which simplifies to \[ \sum_{k=0}^{n}\binom nk=2^n \]

OpenStudy (anonymous):

Remember that: \[ \sum_{k=0}^{n}\binom nk =\binom n0+\binom n1 +\binom n2+\ldots +\binom n{n-1}+\binom nn \]